# Machine learning of the Ising model on a spherical Fibonacci lattice

### v1

@@ III Ferromagnetic Ising model on a spherical Fibonacci lattice
Figure 7: Classification confidences $p_{o}$ and $p_{d}$ versus temperature for the ferromagnetic Ising model on a spherical Fibonacci lattice. The phase transition temperature $T_{c}$ is the point at which $p_{\text{o}}=p_{\text{d}}$.
@@ IV Antiferromagnetic Ising model on a spherical Fibonacci lattice
To determine the phase transition temperature, we will continue using the two previous methods. Following the same procedure, we calculate the total energy for stable spin configurations at each temperature and plot $E$ vs $T$ in Fig.9. By fitting the $E-T$ curve with spine functions and taking derivatives with respect to $T$, we obtain the specific heat $C_{V}$, which is also presented in Fig.9. Interestingly, the total energy near zero temperature is not the lowest, which contrasts with the behavior observed in the planar square lattice. This can be explained as follows. The planar square lattice is a bipartite lattice that can be divided into two sub-lattices, with sites from one sub-lattice interacting only with those from the other. If the spins on the two sub-lattices have opposite signs, the total energy is the lowest. When the spins on the two sub-lattices have opposite signs, the total energy is minimized. In the case of the Fibonacci lattice, the sites with $3$ or $5$ neighbors break this symmetry, leading to an increase in total energy, so that the most uniform spin distribution does not correspond to the lowest energy state. This is, in fact, a distinctive featture of non-uniform lattices. Furthermore, based on the behavior of $C_{V}$, we obtain a rough estimation of the critical temperature: $T_{c}/|J|=2.09$, which is close to the phase transition temperature of the 2D antiferromagnetic Ising model $T^{\Box}_{c}/J\simeq 2.269$[30].
@@
Figure 9: (Left panel) Total energy as a function of temperature for the spherical ferromagnetic Ising model with $N=1000$. (Right panel) Specific heat as a function of temperature for the same model.
The method based on GCN provides a more accurate estimate to the $T_{c}$ of antiferromagnetic Ising model on a spherical Fibonacci lattice. We present our numerical results in Fig.10, where the critical temperature is found to be $T/|J|=2.46$. It is fairly close to $T_{c}$ of the ferromagnetic Ising model on the spherical Fibonacci lattice. This similarity may arise from the resemblance between the spin configurations of the two models at high temperatures: The spin distributions of both are similarly chaotic. Moreover, $p_{\text{o}}$ and $p_{\text{d}}$ exhibit non-smooth behavior in the high-temperature regime. This non-smoothness is due to the inherent ambiguity in the resemblance between spin configurations at low and high temperatures in the spherical antiferromagnetic Ising model, which results in suboptimal performance of the GCN in learning tasks. Nonetheless, despite these challenges, this approach remains instrumental in identifying the phase transition temperature of the model.
@@
Figure 10: Critical temperature of the antiferromagnetic Ising model on a spherical Fibonacci lattice.
@@ V Conclusion
In this paper, we studied the Ising model on a spherical Fibonacci lattice, employing both Monte Carlo simulations and graph convolutional networks(GCNs) to analyze spin configurations and identify phase transition temperatures. Despite the Fibonacci lattice’s relative uniformity, the presence of irregular sites led to significant effects on spin configurations. In the ferromagnetic case, we found that sites with fewer neighbors were more likely to experience spin flips at low temperatures, while this tendency diminished at higher temperatures, contributing to an increased phase transition temperature compared to the planar Ising model. In the antiferromagnetic case, irregularities in the lattice structure resulted in a scenario where the total energy at zero temperature is not at its minimum, reflecting the complexity of interactions influenced by the lattice geometry.
By utilizing specific heat analysis and GCNs, we successfully estimated the phase transition temperatures for both ferromagnetic and antiferromagnetic cases, revealing values that underscore the impact of geometric properties on spin interactions. Our findings highlight the importance of understanding these geometric features in the context of microgravity environments, paving the way for future research in both theoretical and experimental settings. Overall, this work enhances our comprehension of the Ising model in non-planar geometries and opens avenues for further investigation into the behaviors of spin systems in complex lattice structures.
@@ VI Acknowledgments
H.G. was supported by the Innovation Program for Quantum Science and Technology(Grant No.2021ZD0301904) and the National Natural Science Foundation of China(Grant No.12074064). X.-Y. H. was supported by the Jiangsu Funding Program for Excellent Postdoctoral Talent(Grant No.2023ZB611).
Appendix A Methodology of phase classification: Details of GCN

### v2

@@ III.3 Determining $T_{c}$ via the method of machine learning
Figure 9: (Top Panel) Classification confidences, $p_{\text{o}}$ and $p_{\text{d}}$, versus temperature for the ferromagnetic Ising model on a spherical Fibonacci lattice. The phase transition temperature $T_{c}$ is the point at which $p_{\text{o}}=p_{\text{d}}$. (Bottom panel) Error analysis of $T_{c}$ via linear fitting near the point $p_{\text{o}}=p_{\text{d}}$.
To verify the accuracy of the GCN simulations, we first investigate two-dimensional square lattices with sizes of $20\times 20$ and $40\times 40$ as a preliminary test. As illustrated in Fig.8 (with the black solid line representing $p_{\text{o}}$ and the red dashed line representing $p_{\text{d}}$), we observe that as the temperature increases, the black solid line gradually declines from 1, indicating a decreasing similarity between the simulation results and the ordered phase. In contrast, the red dashed line steadily increases from 0, reflecting an increasing similarity between the simulation results and the disordered phase. Through the simulation and analyses of different lattice scales, it is found that as the number of lattice sites increases, the phase transition temperature gradually approaches the theoretical value of $T_{c}/J=2.269$. Specifically, the values obtained are $T_{c}/J=2.51$ for the $20\times 20$ lattice and $T_{c}/J=2.36$ for the $40\times 40$ lattice. The latter demonstrates strong concordance with the predicted critical temperature. This indicates that increasing the lattice size can enhance the accuracy of phase transition temperature estimations. Moreover, the GCN effectively captures the phase transition behavior and critical properties.
After validating the accuracy of GCN, we subsequently apply it to determine the phase transition temperature of the spherical Ising model. In numerical calculations, we select 800 temperature nodes within the range of $[0.01,8.0]J$, with an interval of $0.01J$, and perform 100 simulations for each temperature node. The results, presented in the top panel of Fig.9, indicate that the critical temperature is determined to be $T_{c}/J=2.35$, where $p_{\text{o}}=p_{\text{d}}$. This value is fairly close to those obtained from the analysis of the specific heat $C_{V}$, but it is comparatively higher than that of the square lattice, $T^{\Box}_{c}=2.269J$. Due to the inhomogeneity of the Fibonacci lattice, a complete analytical analysis is not feasible. We believe a possible reason for this is the presence of sites with 3 or 5 neighbors, which affects the spin dynamics within the system. At low temperatures, a site with a greater number of neighbors is less likely to experience a spin flip, as confirmed in Fig.4. However, as the temperature increases beyond $T\gtrsim 2.34J$, the situation changes. Fig.4 shows that at $T=2.34J$, the spin-flip probabilities follow the order $P_{4}>P_{5}>P_{3}$ at $T=2.34J$, indicating that spins on sites with 3 or 5 neighbors are more resistant to flipping. This, to some extent, enhances the “stability” of the ordered phase, effectively elevating the phase transition temperature.
To analyze the uncertainty of $T_{c}$, we note that it is determined by the intersection of the two curves $p_{\text{o}}(T)$ and $p_{\text{d}}(T)$, both of which exhibit good linearity near $T_{c}$. Therefore, the uncertainty can be estimated as follows. We select an interval $[T_{1},T_{2}]$ around $T_{c}$, where $p_{\text{o}}(T)$ and $p_{\text{d}}(T)$ can be approximated as linear functions with slopes $k_{\text{o}}$ and $k_{\text{d}}$, respectively. Each point on these two curves has its own uncertainty, obtained by averaging over 10 stable spin configurations. Using error propagation, we can determine the uncertainties of $k_{\text{o}}$ and $k_{\text{d}}$, denoted as $\sigma_{\text{o}}$ and $\sigma_{\text{d}}$, respectively. Finally, the uncertainty of $T_{c}$ is given by
$\Delta T_{c}=\frac{\sqrt{\sigma_{\text{o}}^{2}+\sigma_{\text{d}}^{2}}}{|k_{\text{o}}-k_{\text{d}}|}.$ (5)
In the bottom panel of Fig.9, we present our numerical results, yielding $T_{c}=2.338\pm 0.0784J$. Note the value $2.338J$ is obtained from the intersection of the two linear fits, which differs slightly from $T_{c}=2.35$, determined by the condition $p_{\text{o}}=p_{\text{d}}$.
III.4 Effects of $r_{c}$ and $N$
$r_{c}/R$ $N_{3}$ ($w_{3}$) $N_{4}$ ($w_{4}$) $N_{5}$ ($w_{5}$) $N_{6}$ ($w_{6}$) $N_{7}$ ($w_{7}$) $N_{8}$ ($w_{8}$)
1.298 76 (7.6%) 850 (85%) 74 (7.4%) - - -
1.500 - 26 (2.6%) 240 (24%) 734 (73.4%) - -
1.700 - - 8 (0.8%) 722 (72.2%) 224 (22.4%) 46 (4.6%)
Figure 10: (Top panel) Classification confidences, $p_{\text{o}}$ and $p_{\text{d}}$, versus temperature for $N=1000$ and different values of $r_{c}$. (Middle panel) Variances of the classification confidences as a function of temperature. (Bottom panel) Percentage of sites with different neighbors for different values of $r_{c}$. Here $N_{i}$ is the number of sites with $i$ neighbors, and $w_{i}=N_{i}/1000$.
Now, we investigate the effects of other parameters, such as $r_{c}$ and $N$, on $T_{c}$. First, we fix the total number of sites at $N=1000$ and vary the interaction length $r_{c}$. As $r_{c}$ increases, each spin interacts with a greater number of spins, effectively increasing the number of neighbors per site and altering the lattice structure. In this scenario, as the temperature rises, the energy cost for a spin flip becomes larger due to the higher average number of neighbors per site, leading to an effective increase in $T_{c}$. We present our numerical results in Fig.10. The top panel shows that $T_{c}=2.35J$ for $r_{c}=0.1298R$, $T_{c}=3.52J$ for $r_{c}=0.1500R$, and $T_{c}=3.95J$ for $r_{c}=0.1700R$, which confirms our former reasoning. Moreover, the error made by GCN when performing phase classification can serve as an indicator for the critical behavior of a dynamical system. Consequently, the temperature at which the classification error is maximized can itself be used as a phase transition criterion [47]. To quantify this, we perform 10 independent simulations at each temperature point and compute the variance of the classification confidence. As shown in the middle panel, the maximum classification errors occur at $T=2.34J$ for $r_{c}=0.1298R$, $T_{c}=3.515J$ for $r_{c}=0.1500R$, and $T_{c}=3.961J$ for $r_{c}=0.1700R$, which agree well with the phase transition temperature obtained from the intersection of the two phases ($p_{\text{o}}$, $p_{\text{d}}$) in classification. This further supports the reliability of the maximum confusion criterion. The bottom panel lists the fraction of sites with different numbers of neighbors for various $r_{c}$ values. For example, when $r_{c}=0.1500R$, $73.4\%$ of the sites have six neighbours, indicating that most of the lattice resembles a triangular structure. Interestingly, the estimated $T_{c}=3.52J$ is relatively close to the exact value $T_{c}=4/\text{ln}3\approx 3.64J$ of the planar triangular lattice as reported in Ref.[48]. Since this model also includes some four-neighbor and five-neighbor sites, its spin flip energy is lower than that of the six-neighbor sites, which results in a relatively lower $T_{c}$. For $r_{c}=0.1700R$, although a similar fraction of sites have six neighbors ($w_{6}=72.2\%$), the number of sites with more neighbors increases significantly ($w_{7}=22.4\%$ and $w_{8}=4.6\%$). This enhanced ”neighbor interaction” effectively strengthens spin correlations, leading to a higher phase transition temperature of $T_{c}=3.92J$.
$N$ $r_{c}$ $N_{3}$ ($w_{3}$) $N_{4}$ ($w_{4}$) $N_{5}$ ($w_{5}$)
500 1.8130 68 (11.6%) 418 (83.6%) 14(2.8%)
1000 1.2980 76 (7.6%) 850 (85%) 74 (7.4%)
3000 0.7470 186 (6.2%) 2740 (91.3%) 74 (2.5%)
5000 0.5818 112 (2.2%) 4740 (94.8%) 148 (3%)
Figure 11: (Top panle) Classification confidences, $p_{\text{o}}$ and $p_{\text{d}}$, versus temperature for $(N,r_{c}/R)=(1000,0.1298)$, (3000,0.747) and (5000,0.5818). (Bottom panel) Corresponding fractions of sites with different numbers of neighbors.
Next, we vary the number of lattice sites and adjust $r_{c}$ accordingly to ensure the majority of sites have four neighbors. However, due to the absence of translational and rotational symmetry in the spherical Fibonacci lattice, this adjustment does not necessarily guarantee that lattices with different site numbers share similar geometric structures. This contrasts with the 2D square lattice, where structural consistency is maintained. In addition to the previously discussed case of $N=1000$, $r_{c}/R=0.1298$, we also select $N=500$, $r_{c}/R=1.8130$; $N=3000$, $r_{c}/R=0.7470$ and $N=5000$, $r_{c}/R=0.5818$, ensuring that the fractions of sites with four neighbors in the latter three cases are 83.6%, 91.3% and 94.8%, respectively. Interestingly, this suggests that the “irregular area” gradually diminishes as the effective curvature of the spherical lattice decreases. The numerical results are visualized in the top panel of Fig.11. Similarly, we determine the phase transition temperatures using two different methods mentioned before. First, by identifying the intersection of the classification confidences, we obtain $T_{c}=2.31J$, $2.35J$, $2.34J$ and $2.33J$ for $N=500$, $1000$, $3000$ and $5000$, respectively. Second, using linear fitting near $T_{c}$ (see the bottom panel of Fig.9), the corresponding phase transition temperatures are giving $T_{c}=2.330\pm 0.1130J$, $2.338\pm 0.0784J$, $2.336\pm 0.0425J$ and $2.347\pm 0.0479J$ for the corresponding system sizes. Surprisingly, at least within the margin of error, these results do not reflect the impact of finite system size on the phase transition temperature. We hypothesize that the following reasons may contribute to this observation: 1. Finite-size effects may not be prominent enough. Although the number of sites increases tenfold from $500$ to $5000$ for the same spherical surface, the interaction range $r_{c}/R$ only decreases from $1.813$ to $0.5818$, a reduction by a factor of $3.12$. More pronounced finite-size effects may require considering systems with significantly larger numbers of sites, which would demand greater computational resources. 2. The folding of the lattice onto the sphere preserves uniformity, maintaining a similarity between local and global behaviors $N\to\infty$. This inherent uniformity may suppress noticeable finite-size effects. All these findings suggest that even for a simple system like the Ising model, the spherical Fibonacci lattice exhibits properties distinct from those of planar lattices, similar to the observations in Ref.[27].
@@ IV Antiferromagnetic Ising model on a spherical Fibonacci lattice
For 2D square lattice, the phase transition temperature of the antiferromagnetic Ising model can also be theoretically predicted, $T^{\Box}_{c}/J\simeq 2.269$[31], which is the same as that of the ferromagnetic Ising model. To determine $T_{c}$ for the spherical Fibonacci lattice, we will continue applying the previous methods. However, due to the less regular spin patterns and the presence of irregular neighbors on the sphere, the antiferromagnetic model exhibits exotic boundaries even at zero temperature. Therefore, the previous methods for estimating entropy and magnetic susceptibility are not applicable in this case. We can only calculate the specific heat by using $C_{V}=\left(\frac{\partial E}{\partial T}\right)_{V}$ and Eq.(3).
@@
Figure 13: (Top panel) Total energy as a function of temperature for the spherical antiferromagnetic Ising model with $N=1000$. (Middle panel) Specific heat as a function of temperature for the same model. (Bottom panel) Specific heat as a function of temperature, calculated using Eq.(3).
Following the same procedure, we compute the total energy for stable spin configurations at each temperature and plot $E$ vs $T$ in Fig.13 (top panel). By fitting the $E-T$ curve with spine functions and taking derivatives with respect to $T$, we obtain the specific heat $C_{V}$, which is also presented in Fig.13 (middle panel). Interestingly, the total energy near zero temperature is not the lowest, which contrasts with the behavior observed in the planar square lattice. This can be explained as follows. The planar square lattice is a bipartite lattice that can be divided into two sub-lattices, with sites from one sub-lattice interacting only with those from the other. If the spins on the two sub-lattices have opposite signs, the total energy is minimized. In the case of the Fibonacci lattice, the sites with 3 or 5 neighbors break this symmetry, leading to an increase in total energy, so that the most uniform spin distribution does not correspond to the lowest energy state. This is, in fact, a distinctive feature of non-uniform lattices. Furthermore, based on the behavior of $C_{V}$, we obtain an estimation of the critical temperature: $T_{c}/|J|=2.07$, which is relatively lower than that of the ferromagnetic case. This is understandable. In an ordered phase at very low temperatures, most sites have 4 neighbors, of which 2 share the same spin. As the temperature increases, it costs $4J$ for such a site to flip its spin, which is only half of the cost in the ferromagnetic Ising model. Consequently, the phase transition temperature is effectively lowered. After performing the Bootstrap error analysis, the result is refined to $T_{c}=2.059\pm 0.016J$.
To cross-check our results, we also employ Eq.(3) to analyze the behavior of $C_{V}$ and present the numerical results in the bottom panel of Fig.13. Notably, the oscillations in this case are more pronounced than those in the ferromagnetic scenario. This is because the spin pattern of the ordered phase is less regular compared to the latter. In the ferromagnetic case, the ordered spin configuration is relatively simple, with most sites sharing the same spin orientation. By fitting the data with a Gaussian function, we estimate the critical temperature as $T_{c}=1.95J$. Furthermore, Bootstrap error analysis refines this estimate to $T_{c}=1.932\pm 0.167J$.
@@
Figure 14: (Top panel) Critical temperature of the antiferromagnetic Ising model on a spherical Fibonacci lattice. (Bottom panel) The corresponding error analysis of $T_{c}$ via linear fitting near the point $p_{\text{o}}=p_{\text{d}}.$
Finally, we apply the method of GCN to estimate the phase transition temperature. The corresponding numerical results are visualized in the top panel of Fig.14, where the critical temperature is found to be $T_{c}/|J|=2.46$. In the bottom panel, we present the linear fittings of $p_{\text{o}}(T)$ and $p_{\text{d}}(T)$ near $T_{c}$. Using Eq. (5), the estimated uncertainty is $T_{c}=2.4895\pm 0.0610J$. The value of $T_{c}$ is fairly close to that obtained for the ferromagnetic case using the same method. This similarity may arise from the resemblance between the spin configurations of the two models at high temperatures, where both exhibit similarly chaotic spin distributions. Furthermore, the estimated $T_{c}$ is also relatively higher than that derived from the behavior of $C_{V}$. This discrepancy could be due to the fact that, in the antiferromagnetic case, the distinction between the ordered and disordered phases near $T_{c}$ is less pronounced than in the ferromagnetic case, which affects the GCN’s ability to differentiate between the two phases. Therefore, we believe that the $T_{c}$ obtained from the former method is relatively more accurate. Additionally, $p_{\text{o}}$ and $p_{\text{d}}$ exhibit non-smooth behavior in the high-temperature regime. This irregularity arises from the inherent ambiguity in the similarity between spin configurations at low and high temperatures in the spherical antiferromagnetic Ising model, leading to suboptimal performance of the GCN in classification tasks. Nevertheless, despite these challenges, the GCN-based approach remains a valuable tool for identifying the phase transition temperature of the model.
@@ V Conclusion
In this paper, we studied the Ising model on a spherical Fibonacci lattice, a structure that balances relative uniformity with irregular sites that significantly influence spin behavior. Employing Monte Carlo simulations and graph convolutional networks (GCNs), we analyzed spin configurations and identified phase transition temperatures for both ferromagnetic and antiferromagnetic cases.
In the ferromagnetic scenario, sites with fewer neighbors were more prone to spin flips at low temperatures, an effect that weakened as temperature rose. This behavior led to a phase transition temperature higher than that of the planar Ising model, highlighting how the lattice’s curvature and connectivity bolster ferromagnetic order. Conversely, in the antiferromagnetic case, lattice irregularities prevented the total energy from reaching its minimum at zero temperature-unlike an ideal bipartite lattice-revealing the complex interplay between geometry and spin interactions in non-planar systems.
By utilizing specific heat analysis and GCNs, which adeptly capture the lattice’s geometric properties, we accurately estimated phase transition temperatures for both cases. Notably, finite size effects were minimal within the system sizes studied, likely due to the uniform distribution of sites in the spherical Fibonacci lattice, mitigating size-dependent variations common in other lattice types. These findings underscore the pivotal role of geometric features in shaping spin dynamics and phase transitions, distinguishing this system from traditional planar models.
This work enhances our understanding of the Ising model in non-planar geometries and holds particular relevance for spin systems in unique settings, such as microgravity environments, where planar assumptions may not apply. By elucidating the effects of curvature and connectivity on spin interactions, our study paves the way for future theoretical and experimental investigations into the behavior of spin systems on complex lattice structures.
@@ VI Acknowledgments
H.G. was supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904) and the National Natural Science Foundation of China (Grant No. 12074064). X. Y. H. was supported by the the National Natural Science Foundation of China (Grant No. 12405008) and the Jiangsu Funding Program for Excellent Postdoctoral Talent (Grant No. 2023ZB611).
Appendix A Bootstrap Error Analysis Method
To assess the statistical uncertainty of the estimation of $T_{c}$ based on the behavior of $C_{V}$, we employ the nonparametric Bootstrap resampling method. This technique simulates statistical fluctuations in the sample distribution without assuming that the data follows a specific distribution, making it well-suited for error estimation in finite-sized systems. Using Eq.(3) as an example, the detailed steps are as follows:
  1. 1.
    Data Preparation
    Suppose the system undergoes Monte Carlo simulations at temperature points $\{T_{i}\}_{i=1}^{N}$. At each temperature point, $M$ independent energy observations are collected, forming the dataset $\mathcal{D}=\{E_{i1},E_{i2},\dots,E_{iM}\}_{i=1}^{N}$.
    $\mathcal{D}=\{E_{i1},E_{i2},\dots,E_{iM}\}_{i=1}^{N}.$
  2. 2.
    Resampling Generation
    For each temperature point $T_{i}$, perform random sampling with replacement:
    $\mathcal{D}^{(b)}_{i}=\left\{E_{ik}^{(b)}\mid k\in\{1,2,\dots,M\},\quad E_{ik}^{(b)}\sim\mathcal{D}_{i}\right\},$
    where the superscript $(b)$ denotes the $b$-th Bootstrap iteration (with a total of $B=1000$ iterations).
  3. 3.
    Statistic Calculation
    For each Bootstrap sample $\mathcal{D}^{(b)}$, compute the target statistics:
    • Specific Heat Capacity:
      $C^{(b)}(T_{i})=\frac{\langle E^{2}\rangle^{(b)}_{i}-\langle E\rangle^{(b)2}_{i}}{T_{i}^{2}}.$
    • Critical Temperature $T_{c}^{(b)}$: Fit the $C^{(b)}(T)$ curve with a Gaussian function and locate the peak temperature.
  4. 4.
    Error Estimation
    • Standard Deviation:
      $\sigma_{T_{c}}=\sqrt{\frac{1}{B-1}\sum_{b=1}^{B}\left(T_{c}^{(b)}-\bar{T}_{c}\right)^{2}},\quad\bar{T}_{c}=\frac{1}{B}\sum_{b=1}^{B}T_{c}^{(b)}.$
    • Confidence Interval: Take the 2.5% and 97.5% percentiles of $\{T_{c}^{(b)}\}$ to define the 95% confidence interval $[T_{c}^{\text{low}},T_{c}^{\text{high}}]$.
Appendix B Method of estimating the entropy
Figure 15: Boundary of a stable spin configuration and its zoom-in view.
To calculate the entropy of the spherical Ising model, we employ the Boltzmann formula:
$S=k_{B}\ln\Omega,$ (6)
where the Boltzmann constant is set to $k_{B}=1$ in natural unites. For a model with $N_{c}$ “proper” sites at the boundaries, the total number of possible states is given by $\Omega=2^{N_{c}}$ since each spin has two orientations. This yields $S=N_{c}\ln 2$. Thus, our primary focus is on determining the “proper” number $N_{c}$ of boundary sites .
Figure 16: A special local structure of the spherical lattice.
The left panel of Fig.15 presents a stable spin configuration, and a zoomed-in view of a part of its boundary is shown in the right panel. Here the boundary refers to the interface between regions of opposite spin orientations, visually represented by the white and black lattice sites. To illustrate the method for determining $N_{c}$, we select a lattice site and identify its nearest neighbor with an opposite spin. For instance, the nearest neighbor of lattice site 2 with an opposite spin is lattice site $B$, and so forth. However, this method results in double-counting of the boundary lattice sites, yielding $N_{c}=6+6=12$. To correct for this, we include the lattice sites themselves in the counting process. Specifically, for lattice site 2, we record its nearest neighbor with opposite spin as site $B$, denoted as $(2,B)$. Similarly, for site 3, we record $(3,C)$, and so on. Likewise, for lattice site $B$, we record $(B,2)$. Since $(2,B)$ and $(B,2)$ are identical, we eliminate one duplicate entry. This refinement leads to a final determination of $N_{c}=6$.
Now consider a more complex configuration. For instance, Fig.16 illustrates the local lattice structure of the spherical Ising model, where a single spin at site 7 is flipped downward. The boundary of this site consists of four neighboring lattice sites $(a,b,c,d)$. To ensure consistency with the condition for determining shown in Fig.15, we also include the lattice site itself as part of the boundary. Consequently, in this case, the ”proper” number of boundary sites is 5: $(7,a,b,c,d)$, leading to $N_{c}=5$.
Appendix C Methodology of phase classification: Details of GCN